Could you please explain to me how to do problems such as 2.4=m+3.7
Answer by Nate(3500)

2.4 = m + 3.7 original
2.4 - 3.7 = m + 3.7 - 3.7 subtract 3.7 to each side in order to have a result for just the variable
-1.3 = m